Circular No. 8993

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 T2 (CARDINAL) An apparently asteroidal object discovered by Rob D. Cardinal, Roth-

ney Astrophysical Observatory, with the University of Calgary 0.50-m f/1.0 reflector at Priddis (discovery observation tabulated below), and posted on the Minor Planet Center’s ‘NEOCP’ webpage has been found to show cometary appearance by other CCD astrometrists. W. H. Ryan (Magdalena Ridge Observatory, Socorro, NM, U.S.A., 2.4-m f/8.9 reflector, Oct. 8.38–8.39 UT, poor seeing) writes that the object appears somewhat fuzzy with an apparent tail in p.a. 315^{o}. G. Hug (Scranton, KS, U.S.A., 0.56-m reflector, Oct. 8.4) finds a short, diffuse coma extension in p.a. ∼ 300^{o}. Y. Ikari (Moriyama, Shiga-ken, Japan, 0.26-m reflector, Oct. 8.4) notes a coma diameter of 0′.3. J. Young (Table Mountain 0.61-m reflector, Oct. 8.5) reports a bright round coma of mag 14.5 and diameter 12′′ with a wide, fan-shaped tail spanning p.a. 240^{o} (where its edge is poorly defined) to 330^{o} (where its edge is well defined, and where it extends to 1′).

2008 UT  α_{2000}  δ_{2000}  Mag.
Oct. 1.16119  7^{h}05^{m}15.^{s}37  +80^{o}03′54′′.8  16.0

The available astrometry, preliminary parabolic orbital elements (T = 2009 June 16.292 TT, q = 1.22108 AU, ω = 214^{o}.987, Ω = 310^{o}.793, i = 56^{o}.124, equinox 2000.0), and an ephemeris appear on MPEC 2008-T88.

COMET P/2008 T3 (BOATTINI) A. Boattini reports his discovery of a comet in a dense part of the

Milky Way on CCD images obtained with the 0.68-m Schmidt telescope in the course of the Catalina Sky Survey (discovery observation tabulated below), with four co-added 30-s exposures showing a strongly condensed 8′′–10′′ coma and a faint, fan-shaped tail extending for ∼ 25′′–30′′ and spanning p.a. 105^{o}–210^{o}; four co-added 30-s exposures by Boattini on Oct. 8.1 UT with the Mt. Lemmon 1.5-m reflector show a coma diameter of 9′′; the brighter tail structure in p.a. 130^{o} extends for ∼ 25′′.

2008 UT  α_{2000}  δ_{2000}  Mag.
Oct. 7.22226  21^{h}42^{m}00.^{s}28  +36^{o}23′55′′.1  17.5

The available astrometry, preliminary elliptical orbital elements (T = 2008 Oct. 25.014 TT, q = 1.14631 AU, e = 0.64958, ω = 181^{o}.311, Ω = 204^{o}.121, i = 33^{o}.040, equinox 2000.0, P = 5.92 years), and an ephemeris appear on MPEC 2008-T89.

2008 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green